# MVSBoost: An Efficient Point Cloud-based 3D Reconstruction

### v1

@@ 2.1 Neural Implicit Fields
In recent years, neural implicit fields have become increasingly popular for multi-view 3D surface reconstruction. The reason for this growth can be attributed to the introduction of differentiable rendering of implicit functions training methods. In early works, surface rendering procedures were relied upon, wherein the color of a pixel was estimated using the radiance of a single point in the volume [Niemeyer2020, 17]. However, these methods have been surpassed by training procedures based on volume rendering, with multiple samples being taken via ray marching.
@@
The main contributions of our paper are as follows: 1. We propose a robust MVS framework:that integrates multi-view $360^{0}$ imagery with advanced processing techniques to produce detailed and accurate 3D point clouds and meshes. 2. We achieve superior performance:outperforming existing neural implicit field approaches and traditional MVS methods in accuracy and precision, as validating through extensive testing with Chamfer distance metrics on a comprehensive dataset. 3. We prove high computational efficiency:by optimizing each stage of the MVS process, our approach substantially reduces the computational resources required, facilitating faster processing and greater efficiency. 4. We illustrate the robustness on complex scene reconstruction:since our method exhibits exceptional capability in handling occlusions and adapting to varying viewpoints, ensuring reliable and comprehensive scene reconstructions. 5. Our method achieves scalability and application versatility:designed to accommodate both bounded and unbounded scenes without specific image-capturing patterns, the proposed framework adapts seamlessly across various scales and environments, making it suitable for various applications.
@@ 3 Proposed Methodology
Our study focuses on achieving precise point cloud-based 3D reconstruction of objects from multiple views.
@@ 3.2 Our Proposal: MVSBoost
Our framework aims to propose rich and expressive meshes from a set of input images =$\mathcal{I}=\{I_{i}|i=1...N_{I}\}$. In this section, we present the Camera Pose and Point Cloud Estimation. Secondly, we present the Densify Point Cloud, Mesh Reconstruction, Mesh Refinement, and Texture Mesh as 3D reconstruction using Point Cloud phases.
@@ 3.3 Camera Pose and Point Cloud Estimation
In Densify Point Cloud, the dense point cloud extraction involves the process of enhancing a sparse point cloud created through SfM by utilizing depth map computation and Depth map fusion techniques. Depth map fusion works by generating depth maps for individual input images through feature matching across adjacent perspectives; the feature matching is conducted at various resolution levels until a detailed set of dense depth maps is obtained. Subsequently, these depth maps are consolidated into a unified dense point cloud, effectively eliminating superfluous points and noise during the fusion process, as shown in Fig. 2 (a).
Mesh reconstruction uses an MVS approach to create a mesh from a dense point cloud. The point cloud is first transformed into a tetrahedral mesh through Delaunay triangulation. Then, a graph-cut optimization determines whether each tetrahedron is inside or outside the object. Finally, the marching cubes algorithm extracts the mesh surface from the labelled tetrahedra. The resulting mesh is a seamless and accurate representation of the object’s geometry based on the original point cloud data, as shown in Fig. 2 (b).
Mesh refinement is a set of techniques used to improve reconstructed meshes’ quality. These techniques include mesh simplification, which reduces the number of vertices while preserving important details, and mesh smoothing, achieved using Laplacian or bilateral filtering to eliminate noise and outliers. Additionally, mesh denoising is used further to enhance the mesh quality through normal voting tensor filtering. Mesh optimization techniques such as vertex relaxation and edge flipping are employed to refine the quality of the triangles. These refinement processes create a clean and precise mesh that accurately represents the object’s surface characteristics, as shown in Fig. 2 (c).
Texture mesh involves mapping images onto the 3D model to create a realistic and detailed surface representation. This process significantly enhances the visual quality of the 3D model, making it suitable for various applications like virtual reality, gaming, and visualization. Texture mapping is crucial for adding color, patterns, and details to the mesh, providing a more immersive and realistic experience when interacting with the 3D model, as shown in Fig. 2 (d).
@@ 4.2 Implementation settings
We used GeForce GTX 1080 Ti/12G to run our experiments. For the Camera Pose and Point Cloud estimation phase (Fig. 1(a)), we use COLMAP [26, 27]. For the feature extractor, we set a single camera option, maximum image size as $1000$, default focal length as $5.2$, maximum number of SIFT features as $2048$, BA global function tolerance as $0.000001$, and focal length ratio as $[0.1\dots 10]$. For the 3D reconstruction phase (Fig. 1(c-f)), we use OpenMVS [28]. We set the close-holes parameter as $400$, mesh smooth parameter as $5$, and the maximum resolution parameter as $512$.
@@ 4.3 Datasets
Our primary experiments use the widely recognized 3D reconstruction benchmark datasets, Realistic Synthetic 360 [18]. The Realistic Synthetic 360 dataset was created for novel view synthesis [18]. Recently, it has been repurposed to evaluate 3D reconstruction algorithms. The dataset contains eight scenes (400 images per scene; each image is 800x800px, including masks and depth images) with complex geometries and non-Lambertian materials, making it a challenging benchmark. The ground truth meshes are filtered to remove non-visible internal surfaces to ensure fair comparisons. Similarly, the reconstructed meshes also undergo a similar process.
@@ 4.4 Comparative Analysis
We compared our method to seven state-of-the-art methods on 8 scenes of the dataset. Each scene presents unique structural complexities in the reconstructed surface, which poses a challenge for standard methods to estimate the locations of high-density regions accurately. This limitation hinders the optimization process and diminishes reconstruction quality and rendering performance.
Quantitative Results: We evaluate the performance of our proposed method with seven state-of-the-art methods (COLMAP [29], NeuS [30], Sphere NeuS [23], NeuS w/ masks [30], Sphere NeuS w/ masks [23], NeuralWarp [4], and Sphere NeuralWarp [23]) on the Realistic Synthetic 360 dataset. The results are shown in Table 1. Our method shows considerable improvements in all scenes compared to the state-of-the-art. Significant enhancements were observed consistently across all Realistic Synthetic 360 dataset scenes.
@@
Qualitative Results: Figure 3 visually compares our method with the ground truth on the Realistic Synthetic 360 dataset. Our method generates reconstructions that are more accurate and closer to the ground truth than the state-of-the-art methods. The Realistic Synthetic 360 dataset provides high-precision ground truth 3D surface models that enable us to quantitatively evaluate the 3D reconstruction algorithms. By visually comparing the reconstructions produced by our method with the ground truth models, we can evaluate the accuracy and completeness of our results. The figure illustrates that our method can closely match the detailed geometry and surface properties of the ground truth, indicating its efficiency in reconstructing 3D objects from multi-view images.
Figure 3: A side-by-side comparison showing our method’s results next to the ground truth GT from the Realistic Synthetic 360 dataset.
@@ 4.5 Discussion
The experimental outcomes validate our hypothesis that traditional MVS techniques when augmented with modern computational methods and camera technologies, can surpass the performance of more recent neural approaches in real-world applications. Integrating SfM algorithms and the strategic use of 360-degree cameras have proven pivotal in enhancing the reconstruction quality. Our results also suggest that there is considerable potential to further optimize these processes, particularly in the aspects of real-time processing and handling dynamic scenes. The experimental results demonstrate that our method outperforms the state-of-the-art methods regarding Chamfer distance on the Realistic Synthetic 360 dataset. The visual results also show that our method can reconstruct more accurate and detailed 3D objects. These results suggest that our method is effective in 3D reconstruction tasks and can be applied to various applications such as robotics, computer vision, and virtual reality. Our results also indicate that there is significant potential to further optimize these processes, particularly in the aspects of real-time processing and handling dynamic scenes.
Complexity: One key advantage of our proposed MVS framework is its computational efficiency. By optimizing various stages of the reconstruction process, such as implementing GPU acceleration and improving the algorithms for image processing and mesh reconstruction, our method significantly reduced the computational time compared to traditional MVS methods. Furthermore, memory consumption was also minimized, enhancing the system’s capacity to handle larger datasets effectively.
Corner cases: Handling of Complex Scenes and Occlusions. The enhanced MVS method performed better in reconstructing complex scenes, particularly those with significant occlusions and diverse viewpoints. The robustness of our approach was evident as it consistently maintained high levels of detail and accuracy in areas where previous methods struggled. This success is attributed to the comprehensive scene coverage provided by 360-degree imagery and the refined techniques for camera pose estimation and point cloud interpolation.
Limitations. The limitations of our method are presented, respectively: 1. Our method requires a high amount of images to have a perfect reconstruction. 2. Our method might not have a good construction for high-frequency areas such as reflections. 3. Our method method may not work for different lighting conditions in the scene.
@@ 5 Conclusions and Future Work
This study successfully demonstrates the enhanced capabilities and significant advantages of our proposed MVS framework over existing neural implicit field methods. Through rigorous experimentation and analysis using the Realistic Synthetic 360 dataset and Chamfer distance metrics, our method was proven to be more accurate, precise, and efficient regarding computational resources. The novel integration of multi-view 360-degree imagery with advanced SfM techniques, alongside subsequent point cloud densification, mesh reconstruction, and texturing processes, ensures the production of high-quality 3D models. These models exhibit enhanced clarity and detail, showcasing the robustness of our approach in handling complex scenes, occlusions, and diverse viewpoints effectively. Our findings underscore the practicality of MVS in real-time and large-scale applications, suggesting that it remains a vital technology in 3D reconstruction, particularly suitable for AR/VR, medical imaging, and media production. Future work will focus on refining the MVS process, exploring more efficient algorithms for image processing, and extending the applicability to even more challenging environments. By continuing to advance the capabilities of MVS, we aim to maintain its relevance and superiority in the evolving landscape of 3D reconstruction technologies.

### v2

@@ 2.1 Neural Implicit Fields
In recent years, neural implicit fields have become increasingly popular for multi-view 3D surface reconstruction. The reason for this growth can be attributed to the introduction of differentiable rendering of implicit functions training methods. In early works, surface rendering procedures were relied upon, wherein the color of a pixel was estimated using the radiance of a single point in the volume [17, 18]. However, these methods have been surpassed by training procedures based on volume rendering, with multiple samples being taken via ray marching.
@@
The main contributions of our paper are as follows: 1. We propose a robust MVS frameworkthat integrates multi-view $360^{0}$ imagery with advanced processing techniques to produce detailed and accurate 3D point clouds and meshes; 2. We achieve superior performance in terms of accuracy and precision, outperforming existing neural implicit field approaches and traditional MVS methods, as demonstrated through extensive testing with Chamfer distance metrics on a comprehensive dataset. 3. We demonstrate a high computational efficiencyby optimizing each stage of the MVS process, leading to a substantial reduction of the computational resources required, facilitating faster processing and greater efficiency; 4. We illustrate the robustness of our approach on complex scene reconstruction: since our method exhibits exceptional capability in handling occlusions and adapting to varying viewpoints, ensuring reliable and comprehensive scene reconstructions. 5. Our method achieves scalability and application versatility:designed to accommodate both bounded and unbounded scenes without specific image-capturing patterns, the proposed framework adapts seamlessly across various scales and environments, making it suitable for various applications.
@@ 3 Proposed Methodology
Our study focuses on achieving precise point cloud-based 3D reconstruction of objects from multiple views. This section introduces our novel pipeline, which embodies our key contributions to the field. Central to our approach is the innovative integration of camera pose estimation (SfM) with subsequent stages: Densify Point Cloud, Mesh Reconstruction, Mesh Refinement, and Texture Mesh. By meticulously coordinating these modules, we have achieved superior performance in terms of accuracy and precision compared to existing state-of-the-art methods. Our pipeline not only enhances computational efficiency but also demonstrates robust capability in handling complex scene reconstruction.
@@ 3.2 Our Proposal: MVSBoost
Our framework aims to propose rich and expressive meshes from a set of $N_{I}$ input images =$\mathcal{I}=\{I_{i}|i=1\ldots N_{I}\}$. In this section, we present the Camera Pose and Point Cloud Estimation. Secondly, we present the Densify Point Cloud, Mesh Reconstruction, Mesh Refinement, and Texture Mesh as 3D reconstruction using Point Cloud phases.
@@ 3.3 Camera Pose and Point Cloud Estimation
In Densify Point Cloud, the dense point cloud extraction involves the process of enhancing a sparse point cloud created through SfM by utilizing depth map computation and Depth map fusion techniques. Depth map fusion works by generating depth maps for individual input images through feature matching across adjacent perspectives; the feature matching is conducted at various resolution levels until a detailed set of dense depth maps is obtained. Let $\mathrm{D}_{i}$ represent the depth map for image $I_{i}$, computed through feature matching: $\mathrm{D}_{i}=f(I_{i},I_{adj},R)$. Where $I_{adj}$ denotes the set of adjacent images and $R$ represents the resolution and different levels of detail. Subsequently, these depth maps are consolidated into a unified dense point cloud $P_{d}$, effectively eliminating superfluous points and noise during the fusion process: $P_{d}=g(\mathrm{D}_{i})$, as shown in Fig. 2 (a).
Mesh reconstruction uses an MVS approach to create a mesh from a dense point cloud. The point cloud is first transformed into a tetrahedral mesh $T$ through $f^{D}$ Delaunay triangulation function: $T=f^{D}(P_{d})$. Then, a graph-cut optimization $f^{G}$ determines whether each tetrahedron is inside or outside the object: $L=f^{G}(T)$. Finally, the marching cubes algorithm $\mathbb{M}$ extracts the mesh surface from the labeled tetrahedra: $M=\mathbb{M}(T,L)$. The resulting mesh $M$ is a seamless and accurate representation of the object’s geometry based on the original point cloud data, as shown in Fig. 2 (b).
Mesh refinement is a set of techniques used to improve the quality of reconstructed meshes. These techniques include mesh simplification $f^{S}$, which reduces the number of vertices while preserving important details: $M_{s}=f^{S}(M)$, and mesh smoothing $f^{S}_{m}$, achieved using Laplacian or bilateral filtering to eliminate noise and outliers: $M_{sm}=f^{S}_{m}(M_{s})$. Additionally, mesh denoising $f^{D}_{e}$ is used further to enhance the mesh quality through normal voting tensor filtering: $M_{d}=f^{D}_{e}(M_{sm})$. Mesh optimization techniques $\hat{f}$, such as vertex relaxation and edge flipping, are employed to refine the quality of the triangles: $\hat{M}=\hat{f}(M_{d})$. These refinement processes create a clean and precise mesh $\hat{M}$ that accurately represents the object’s surface characteristics, as shown in Fig. 2 (c).
Texture mesh involves mapping images onto the 3D model to create a realistic and detailed surface representation. This process significantly enhances the visual quality of the 3D model, making it suitable for various applications like virtual reality, gaming, and visualization. Texture mapping involves associating 2D image coordinates (texture coordinates) with the vertices of the 3D mesh: $T_{c}=\{(u_{i},v_{i})|i=1,...,N_{v}\}$. Apply the texture coordinates $f^{T}$ to the mesh $\hat{M}$ to obtain the textured mesh: $M_{t}=f^{T}(\hat{M},T_{c},\mathcal{I})$. Texture mapping is crucial for adding color, patterns, and details to the mesh, providing a more immersive and realistic experience when interacting with the 3D model, as shown in Fig. 2 (d).
@@ 4.2 Implementation settings
We used GeForce GTX 1080 Ti/12G to run our experiments. For the Camera Pose and Point Cloud estimation phase (Fig. 1(a)), we use COLMAP [27, 28]. For the feature extractor, we set a single camera option, maximum image size as $1000$, default focal length as $5.2$, maximum number of SIFT features as $2048$, BA global function tolerance as $0.000001$, and focal length ratio in range $[0.1\dots 10]$. For the 3D reconstruction phase (Fig. 1(c-f)), we use OpenMVS [29]. We set the close-holes parameter as $400$, mesh smooth parameter as $5$, and the maximum resolution parameter as $512$.
@@ 4.3 Datasets
Our primary experiments use the widely recognized 3D reconstruction benchmark datasets, Realistic Synthetic 360 [19]. The Realistic Synthetic 360 dataset was created for novel view synthesis [19]. Recently, it has been repurposed to evaluate 3D reconstruction algorithms. The dataset contains eight scenes (400 images per scene; each image is $800\times 800$ px, including masks and depth images) with complex geometries and non-Lambertian materials, making it a challenging benchmark. The ground truth meshes are filtered to remove non-visible internal surfaces to ensure fair comparisons. Similarly, the reconstructed meshes also undergo a similar process.
@@ 4.4 Comparative Analysis
We compared our method to seven state-of-the-art methods on eight scenes of the Realistic Synthetic 360 dataset, each presenting unique structural complexities in the reconstructed surface. These complexities challenge standard methods in accurately estimating the locations of high-density regions. This limitation hinders the optimization process and diminishes reconstruction quality and rendering performance.
Quantitative Results: Our proposed method was evaluated against seven state-of-the-art methods (COLMAP [30], NeuS [31], Sphere NeuS [24], NeuS w/ masks [31], Sphere NeuS w/ masks [24], NeuralWarp [4], and Sphere NeuralWarp [24]) using the Realistic Synthetic 360 dataset. The results are shown in Table 1, indicating that our method significantly outperforms all scenes compared to the state-of-the-art.
@@
Figure 3: A side-by-side comparison of our method’s results with the ground truth from the Realistic Synthetic 360 dataset. The figure illustrates significant differences and improvements in the presented scenes: the Chair, Hotdog, and Ship. Our reconstruction exhibits superior precision, refinement, detail, and texture fidelity compared to existing methods, as evidenced by the clear visual distinctions.
Qualitative Results: Fig. 3 compares our method with the ground truth on the Realistic Synthetic 360 dataset. It demonstrates that our reconstructions are more accurate and closer to the ground truth than state-of-the-art methods. The dataset provides high-precision 3D surface models, allowing us to evaluate 3D reconstruction algorithms quantitatively. By comparing our reconstructions with the ground truth, we can assess the accuracy and completeness of our results. The figure shows that our method closely matches the ground truth’s detailed geometry and surface properties, highlighting its efficiency in reconstructing 3D objects from multi-view images.
@@ 4.5 Discussion
The results validate our hypothesis that augmenting traditional MVS techniques with modern computational methods and 360-degree camera technologies can surpass recent neural approaches in real-world applications. Our integration of SfM algorithms and 360-degree cameras enhances reconstruction quality, with substantial potential for further optimization in real-time processing and dynamic scene handling.
Complexity: Our MVS framework optimizes the reconstruction process through GPU acceleration and improved algorithms for image processing and mesh reconstruction, significantly reducing computational time and memory consumption compared to traditional state-of-the-art methods. We re-trained the state-of-the-art approaches on their defined settings using the same machine for a fair comparison. The state-of-the-art methods of the neural implicit field with which we are comparing take considerably longer: Sphere Neus [24] takes 4 hours, 17 minutes, and 26.411 seconds, and Sphere NeuralWarp [24] takes 9 hours, 29 minutes, and 47.156 seconds. In contrast, our proposed method takes only around 9 minutes and 19.85 seconds. This substantial reduction in computational time proves our claim that our proposed method significantly improves efficiency and minimizes memory consumption, thereby enhancing the system’s capacity to handle larger datasets effectively.
Figure 4: A detailed comparison showcasing selected mesh parts from ground truth, our method, and previous state-of-the-art methods, demonstrating robustness by preserving a high level of details in challenging areas where previous methods struggle, as highlighted in this figure.
Corner cases: Our method excels in reconstructing complex scenes with significant occlusions and diverse viewpoints. The robustness of our approach is evident in its consistent detail and accuracy, attributed to comprehensive scene coverage from 360-degree imagery and refined camera pose estimation techniques, as shown in Fig. 4.
Limitations: Despite its advantages, our method has some limitations: 1. Requires a high number of images for perfect reconstruction. 2. Struggles with high-frequency areas such as reflections. 3. May not perform well under varying lighting conditions.
@@ 5 Conclusions and Future Work
This study demonstrates the significant advantages of our proposed MVS framework over existing neural implicit field methods. Using the Realistic Synthetic 360 dataset and Chamfer distance metrics, our method proves to be more accurate, precise, and computationally efficient. The integration of multi-view 360-degree imagery with advanced SfM techniques, point cloud densification, mesh reconstruction, and texturing processes results in high-quality 3D models that effectively handle complex scenes, occlusions, and diverse viewpoints. Our findings highlight the practicality of MVS in real-time and large-scale applications, making it suitable for AR/VR, medical imaging, and media production. Future work will aim to refine the MVS process, explore more efficient algorithms for image processing, and extend applicability to more challenging environments, maintaining its relevance and superiority in 3D reconstruction technologies.